OpenStudy (anonymous):

Identify the 27th term of an arithmetic sequence where a1 = 38 and a17 = -74

OpenStudy (mathstudent55):

An arithmetic sequence has a constant difference between two consecutive terms. If the fisrt term is a1, and the nth term is an, then a1 = a1 a2 = a1 + d a3 = a2 + d = a1 + d + d = a1 + 2d a4 = a3 + d = a1 + 2d + d = a1 + 3d ... an = a1 + (n - 1)d You are given a1 and a17, and for a17, n = 17. Using the formula above, find d. Then use the formula again using the d you found and n=27 to find a27.

OpenStudy (anonymous):

@mathstudent55 I need help too!